two angles of a triangle have the same measure. The measure of the third angle is five less than three times the measure of each of the other angels. Find the measures of the angles in the triangle.
Answer by checkley79(3341)

X+X+(3X-5)=180
5X=180+5
5X=185
X=185/5
X=35 ANS. FOR THE TWO SMAE ANGLES.
3*35-5=105-5=100 ANS. FOR THE THIRD ANGLE.
PROOF:
35+35+100=180
180=180